# o1-Coder: an o1 Replication for Coding

### v1

@@ Abstract
The technical report introduces O1-CODER, an attempt to replicate OpenAI’s o1 model with a focus on coding tasks. It integrates reinforcement learning (RL) and Monte Carlo Tree Search (MCTS) to enhance the model’s System-2 thinking capabilities. The framework includes training a Test Case Generator (TCG) for standardized code testing, using MCTS to generate code data with reasoning processes, and iteratively fine-tuning the policy model to initially produce pseudocode, followed by the generation of the full code. The report also addresses the opportunities and challenges in deploying o1-like models in real-world applications, suggesting transitioning to the System-2 paradigm and highlighting the imperative for environment state updates. Updated model progress and experimental results will be reported in subsequent versions. All source code, curated datasets, as well as the derived models will be disclosed at https://github.com/ADaM-BJTU/O1-CODER .
@@ 1 Introduction
OpenAI recently introduced the o1 model (OpenAI, 2024), which has demonstrated impressive system-2 thinking capabilities. This model represents a significant advancement in AI’s ability to perform complex reasoning tasks that require higher-order cognitive functions. Following its release, numerous analysis and replication efforts have emerged, highlighting the growing interest and potential of o1-like models. Notable works include g1 (Benjamin Klieger, 2024), OpenO1 (ope, 2024), O1-Journey (GAIR-NLP, 2024), OpenR (Team, 2024), LLaMA-O1 (SimpleBerry, 2024), LLaMA-Berry (Zhang et al., 2024), Steiner (Ji, 2024), Thinking Claude (Richards Tu, 2024), LLaVA-o1 (Xu et al., 2024), and several industrial releases such as k0-math, DeepSeek-R1-Lite, Macro-o1 (Zhao et al., 2024), Skywork o1, QwQ (Qwen Team, 2024), and InternThinker (Shanghai AI Lab, 2024) (illustrated in Fig. 1).
@@
Prior to the o1 model, large language models (LLMs) primarily exhibited System-1 capabilities, characterized by fast, intuitive responses. These models were trained on datasets consisting mainly of question-answer $(Q,A)$ pairs, lacking the intermediate reasoning steps that involve deliberate and analytical processing. This stems from the fact that humans rarely record their thought processes on the internet or elsewhere. Traditionally, techniques such as Chain-of-Thought (CoT) prompting were used to guide models in generating step-by-step reasoning before arriving at an answer. However, a more direct and effective way is to create datasets including the reasoning sequences, e.g., $(Q,...,S_{i},...,A)$, where $S_{i}$ represents an individual reasoning step leading to the final answer.
@@
It is widely believed that o1 addresses the lack of reasoning data by combining reinforcement learning with pretraining. Reinforcement learning (RL) is well known for its ability to explore and discover new strategies rather than relying on predefined data. Looking back at key developments in machine learning, we can see that deep learning and large-scale pretraining have driven transformations in model architecture and the requirements for labeled data, respectively. In contrast, reinforcement learning addresses a different aspect of transformation on the objective function. In situations where explicit guidance or clear goals are absent, RL exploits exploration to search for new knowledge and solutions. Combining pretraining with RL creates a powerful synergy of learning and search, where pretraining compresses existing human knowledge, and RL enables the model to explore new possibilities.
@@
We chose coding tasks to explore how to employ RL to generate and refine reasoning data. Coding is a typical task that requires System-2 thinking, involving careful, logical, and step-by-step problem-solving. Moreover, coding can serve as a foundational skill for solving many other complex problems. This technical report presents our attempt to replicate o1 with a specific focus on coding tasks. The approach integrates RL and Monte Carlo Tree Search (MCTS) to enable self-play, allowing the model to continually generate reasoning data and enhance its System-2 capabilities.
@@ 2 Framework Overview
There are two main challenges to address for self-play RL applied to code generation. The first challenge is result evaluation, i.e., assessing the quality of the generated code. Unlike tasks such as Go or mathematics, where results can be directly evaluated based on game rules or correct answers, evaluating code requires running the generated code within a testing environment and verifying it against test cases. We cannot assume that code datasets will always provide sufficient test cases. The second challenge involves defining the thinking and search behaviors, i.e., determining the object and granularity of process rewards. For code generation, the key question is how to design the reasoning process and the space of policies to guide the model’s behavior effectively.
@@
To address the first challenge, we propose training a Test Case Generator (TCG), which automatically generates test cases based on the question and the ground-truth code  [^1] We also propose an alternative approach where test cases are generated based solely on the question. In addition to utilizing code datasets only provide questions, it can also be applied during the inference phase, enabling online reasoning without the need for predefined ground-truth code.. This approach will help build a standardized code testing environment, providing result rewards for reinforcement learning.
@@
For the second challenge, two possible approaches can be considered. One is “think before acting”, where the model first forms a complete chain of thought and then generates the final answer all at once. The other approach, “think while acting” (Zelikman et al., 2024), involves generating parts of the answer while simultaneously reasoning through the task. We chose the former approach. For code generation, this means first thinking through and writing out a detailed pseudocode, which is then used to generate the final executable code. The advantages are twofold: adaptability, as the same pseudocode can lead to different concrete code implementations; and controllable granularity, as adjusting the level of detail in the pseudocode can be adjusted to control the granularity of the reasoning/search behavior.
@@
The complete framework pseudocode is provided in Algorithm 1, which consists of six steps. (1) The first step is training the test case generator (TCG) $\gamma_{\text{TCG}}$, which is responsible for automatically generating test cases based on the question. (2) In the second step, we run MCTS on the original code dataset to generate code data with reasoning processes $\mathcal{D}_{\text{process}}$, including a validity indicator to distinguish between correct and incorrect reasoning steps. (3) Once we have data that includes the reasoning process, the third step is to fine-tune the policy model $\pi_{\theta}$, training it to behave in a “think before acting” manner. (4) The reasoning process data can also be used to initialize the process reward model (PRM) $\rho_{\text{PRM}}$, which evaluates the quality of reasoning steps. (5) The fifth step is the most crucial: with PRM $\rho_{\text{PRM}}$ providing process rewards and TCG $\gamma_{\text{TCG}}$ provides result rewards, the policy model $\pi_{\theta}$ is updated with reinforcement learning and MCTS. (6) In the 6th step, based on the updated policy model, new reasoning data can be generated. This new data can then be used to fine-tune the PRM again (4th step). Therefore, steps 4, 5, and 6 form an iterative cycle, where self-play continues to drive model improvements. The flow between the six steps is illustrated in Fig. 2. The following section will introduce each step in detail.
@@ 3.1.3 Experiments
In practical scenarios, the generator’s performance has generally met the requirements for assessing code correctness. Looking ahead, we plan to explore the DPO method, where the model autonomously generates data during inference, potentially optimizing the generator’s performance further.
@@ 3.2.1 Pseudocode-based Reasoning Process
The definition of the reasoning process is crucial. As mentioned in the Introduction, we explore a pseudocode-based prompting approach designed to guide large language models in deep reasoning for complex code tasks. Pseudocode, serving as an intermediate representation between natural language descriptions and actual code, offers a more abstract and concise way to express the logical flow of algorithms or programs. To integrate pseudocode reasoning into step-level Chain-of-Thought (CoT), as illustrated in Fig. 4, we define three key behavioral actions infused with pseudocode reasoning:
@@ 3.4 PRM Training
Point-wise In this format, data collected from the search tree are organized as $D=\{(Q_{i},S_{i}^{1:j-1},S_{i}^{j},v_{i}^{j})\mid i=1,2,\ldots,N\}$, where $N$ is the number of samples, and $v_{i}^{j}$ represents the value label assigned to step $S_{i}^{j}$ during the tree search process. Depending on the processing method, this label can be used to derive either hard or soft estimates. Following the approach in (Wang et al., 2024), the PRM is trained using the objective:
@@ 4.1 Bitter Lesson: Data is All You Need
Over the last decade, the AI field has been developing along a central line towards maximizing computation-intelligence conversion efficiency, which is to efficiently convert the ever-increasing computing power into higher intelligence levels. Along this line, as illustrated at the top of Fig. 6, early advancements prioritized improvements on the model side: from SVM to DNN and then to Transformer, scalable model architectures were designed to fully leverage computational power.
In recent years, the focus has shifted towards the data side. Techniques such as Semi-Supervised Learning (SSL) in pre-training and Reinforcement Learning (RL) in post-training have aimed to harness data more effectively. The o1 model continues this line. It moves from SFT, which leverages high-quality supervised data, to RLHF, which utilizes environmental feedback to access theoretically unlimited data, and finally to o1’s innovative approach of supervising the generation process through reward signals derived from the generated reasoning process itself.
This progression suggests that, with Transformer architectures now capable of scaling to handle vast amounts of data and training models of sufficient size, the only remaining challenge converges to acquiring adequate data. One approach is to seek data wherever it is lacking, such as reasoning data for system-2 abilities or physical world trajectories for embodied intelligence. Another approach is to explore data types that do not yet exist in the human world, which requires further exploration of techniques like RL and Self-Play.
@@ 4.2 Sweet Lesson: Beyond Human Data
A common criticism of LLM is its reliance on existing human-recorded data, which inherently limits their potential. As Wittgenstein stated, “The limits of my language mean the limits of my world.” The finite scope and depth of human language records constrain the cognitive capabilities of LLMs. However, the success of o1 demonstrates that we can now explore the underlying thought processes behind these recorded data through RL. This advancement signifies a pivotal shift in AI development, moving from mere imitation of human language to the autonomous generation of novel cognitive processes.
More interestingly, these thought process data do not necessarily be confined to natural language. As highlighted in a recent Nature paper (Fedorenko et al., 2024), “language serves primarily as a tool for communication rather than the essence of thought.” In our observations, some of the thought chains generated by o1 contain nonsensical text, suggesting that the thinking tokens may not correspond to discrete natural language words. If the model has developed itself a more efficient form of internal representation for thinking, this will significantly elevate the efficiency of thought processes and problem-solving mechanisms, not only transcending the limitations imposed by human language data but also further unlocking the potential of model capabilities.
4.3 Opportunities
The self-play+RL framework provides a viable solution for exploring underlying data, which opens up the possibility of exploring System-2 solutions for many tasks that were previously reliant on System 1 capabilities. By integrating more thoughtful, step-by-step processes into task execution, we believe that this approach can yield positive results across a wide range of domains (Kant et al., 2024; Ganapini et al., 2021; Valmeekam et al., 2024; Lowe, 2024). Tasks traditionally solved using System 1 capabilities, such as reward modeling (Mahan et al., 2024), machine translation (Zhao et al., 2024), retrieval-augmented generation (RAG) (Li et al., 2024), and multimodal QA (Islam et al., 2024), have already benefited from the deeper reasoning capabilities enabled by System-2 thinking.
The o1 model’s system card demonstrates notable improvements in model safety. Inspired by this, we have recently explored the concept of System-2 Alignment, which involves guiding models to thoroughly evaluate inputs, consider potential risks, and correct biases in their reasoning (Wang & Sang, 2024). We introduced three methods to realize System-2 alignment: prompt engineering, supervised fine-tuning, and reinforcement learning with process supervision. We will apply the Self-Play+RL framework presented in this report to System-2 alignment, aiming to further enhance the model’s ability to think deliberately and reduce vulnerabilities in complex scenarios.
4.4 Challenges
@@
For o1-like inference models to be deployed across broader real-world applications, two major challenges need to be addressed, both involving with the RL environments. The first challenge concerns reward function generalization. This has been already discussed in the community. For example, leveraging the enhanced ability of inference models to understand high-level natural instructions, approaches like Constitutional AI (Bai et al., 2022) might directly define reward functions in natural language. Another strategy focuses on improving coding capability and transforming the other tasks into coding problems for resolution.
@@
Another less mentioned challenge concerns environment state update. Unlike classic model-free RL methods, such as Q-learning, where state transitions are not explicitly modeled, o1-like models rely on behavior simulation and forward search, requiring knowledge of the updated state following an action. This shifts the paradigm towards model-based RL. In well-defined domains such as programming, mathematics, and Go, the environment often has deterministic rules. For example, programming uses compiler-defined language specifications, mathematics adheres to axiomatic logic, and Go operates under fixed game rules. These deterministic frameworks allow precise computation of state transition probabilities $p(\text{state}_{i+1}\mid\text{state}_{i},\text{action}_{i})$ following specific actions.
@@
However, in many real-world applications, such as Retrieval-Augmented Generation (RAG), device usage , and embodied agents, obtaining state updates requires interaction with external environments or simulators. This introduces significant computational and time costs. For example, in device use, behaviors like clicking, inputting, or scrolling must be simulated in a way that involves page rendering, state updates, and sometimes complex backend interactions like network requests. Moreover, o1-like models face the limitation of not being able to perform online behavior simulation during inference, which prevents the model from validating or correcting its actions by returning to a previous state. This leads to inability to backtrack and refine decisions.
@@
Therefore, one of the key directions is to attempt explicit modeling of the environment by developing a world model for state transition prediction. The world model takes as input the current and past states and actions, and produces the next state as output. This allows the model to interact with its internal world model, rather than directly with the real environment or a simulator. We recognize that one of the ongoing challenges in RL when building such world models is ensuring their accuracy. As a result, world models have typically been applied to environments where the dynamics are relatively simple and well-understood. The good news is, recent rapid advancements in generative games (Sang, 2024) offer promising progress that could facilitate more accurate and practical environment modeling for inference models in real-world applications.
@@
Prospects.   The o1 model is clearly influenced by AlphaGo: AlphaGo utilized imitation learning to initialize the policy network, reinforcement learning to fine-tune the policy and learn the value network, and MCTS as an online search strategy, which parallels LLM’s pre-training, post-training, and inference. AlphaGoZero took a more advanced approach by not relying on historical data, which exactly mirrors current trends in LLM development increasingly emphasizing the post-training stage. If we follow the evolution of the Alpha series, we can anticipate similar developments in o1-like inference models. Initially, the Alpha series developed towards generalization: AlphaZero was applied to Go, Chess, and Shogi, while MuZero achieved human-level performance across 57 Atari games. The other goal besides generalization, however, is to apply these models to more complex, real-world tasks. This progression is evident in AlphaFold’s leap to AlphaCode and AlphaGeometry, as well as the extension of AI to physical environments, such as the 3D virtual agents in SIMA or the embodied intelligence in RT-X.

### v2

@@ Abstract
The technical report introduces O1-CODER, an attempt to replicate OpenAI’s o1 model with a focus on coding tasks. It integrates reinforcement learning (RL) and Monte Carlo Tree Search (MCTS) to enhance the model’s System-2 thinking capabilities. The framework includes training a Test Case Generator (TCG) for standardized code testing, using MCTS to generate code data with reasoning processes, and iteratively fine-tuning the policy model to initially produce pseudocode and then generate the full code. The report also addresses the opportunities and challenges in deploying o1-like models in real-world applications, suggesting transitioning to the System-2 paradigm and highlighting the imperative for world model construction. Updated model progress and experimental results will be reported in subsequent versions. All source code, curated datasets, as well as the derived models are disclosed at https://github.com/ADaM-BJTU/O1-CODER.
@@ 1 Introduction
OpenAI recently introduced the o1 model (OpenAI, 2024), which has demonstrated impressive system-2 thinking capabilities. This model represents a significant advancement in AI’s ability to perform complex reasoning tasks that require higher-order cognitive functions. Following its release, numerous analysis and replication efforts have emerged, demonstrating the growing interest in reasoning models. Notable works include g1 (Benjamin Klieger, 2024), OpenO1 (ope, 2024), O1-Journey (GAIR-NLP, 2024), OpenR (Team, 2024), LLaMA-O1 (SimpleBerry, 2024), LLaMA-Berry (Zhang et al., 2024), Steiner (Ji, 2024), Thinking Claude (Richards Tu, 2024), LLaVA-o1 (Xu et al., 2024), and several industrial releases such as k0-math, DeepSeek-R1-Lite, Macro-o1 (Zhao et al., 2024), Skywork o1, QwQ (Qwen Team, 2024), and InternThinker (Shanghai AI Lab, 2024) (illustrated in Fig. 1).
@@
Prior to the o1 model, large language models (LLMs) primarily exhibited System-1 capabilities, characterized by fast, intuitive responses. These models were trained on datasets consisting mainly of question-answer $(Q,A)$ pairs, lacking the intermediate reasoning steps that involve deliberate and analytical processing. This stems from the fact that humans rarely record their thought processes on the internet or elsewhere. Traditionally, techniques such as Chain-of-Thought (CoT) prompting were used to guide models in generating step-by-step reasoning before arriving at an answer. A more direct and effective way is to create datasets including the reasoning sequences, e.g., $(Q,...,S_{i},...,A)$, where $S_{i}$ represents an individual reasoning step leading to the final answer. Thus, a second approach to enhancing System-2 reasoning capabilities is supervised fine-tuning (SFT). However, most publicly available data is recorded in a question-answer form, and annotating or distilling such data, especially for complex tasks, is both costly and challenging. In this study, we aim to explore in the absence of reasoning process data, and thus opt for the third approach of reinforcement learning (RL).
@@
It is widely believed that o1 addresses the lack of reasoning data by combining reinforcement learning with pretraining. Reinforcement learning is well known for its ability to explore and discover new strategies rather than relying on predefined data in the past decade. Looking back at key developments in machine learning, we can see that deep learning and large-scale pretraining have driven transformations in model architecture and the requirements for labeled data, respectively. In contrast, reinforcement learning addresses a different aspect of transformation on the objective function. In situations where explicit guidance or clear goals are absent, RL exploits exploration to search for new knowledge and solutions. Therefore, combining pretraining with RL creates a powerful synergy of learning and search, where pretraining compresses existing human knowledge, and RL enables the model to explore new possibilities.
@@
We chose coding tasks to explore how to employ RL to generate and refine reasoning data. Coding is a typical task that requires System-2 thinking, involving careful, logical, and step-by-step problem-solving. Moreover, coding can serve as a foundational skill for solving many other complex problems. This report presents our attempt to replicate o1 with a specific focus on coding tasks. The approach integrates RL and Monte Carlo Tree Search (MCTS) to enable self-play, allowing the model to continually generate reasoning data and enhance its System-2 capabilities.
@@ 2 Framework Overview
There are two main challenges to address for applying self-play RL to code generation. The first challenge is result evaluation, i.e., assessing the quality of the generated code. Unlike tasks such as Go or mathematics, where results can be directly evaluated based on game rules or correct answers, evaluating code requires running the generated code within a testing environment and verifying it against test cases. We cannot assume that code datasets will always provide sufficient test cases. The second challenge involves defining the thinking and search behaviors, i.e., determining the state transition and the granularity of process rewards. For code generation, the key question is how to design the reasoning process and the space of policies to guide the model’s behavior effectively.
@@
To address the first challenge, we propose training a Test Case Generator (TCG), which automatically generates test cases based on the question and the ground-truth code  [^1] We also propose an alternative approach where test cases are generated based solely on the question. In addition to be capable of utilizing code datasets that only provide questions, it can also be applied during the inference phase, enabling online reasoning without the need for ground-truth code.. This approach will help build a standardized code testing environment, providing result rewards for reinforcement learning.
@@
For the second challenge, two possible approaches can be considered. One is “think before acting”, where the model first forms a complete chain-of-thought and then generates the final answer all at once. The other approach, “think while acting” (Zelikman et al., 2024), involves generating parts of the answer while simultaneously reasoning through the task. We chose the former approach in this study. For code generation, this means first thinking through and writing out a detailed pseudocode, which is then used to generate the final executable code. The advantages are two-fold: adaptability, as the same pseudocode can lead to different concrete code implementations; and controllable granularity, as adjusting the level of detail in the pseudocode can control the granularity of the reasoning/search behavior.
@@
The outlined framework is provided in Algorithm 1, which consists of six steps. (1) The first step is training the test case generator (TCG) $\gamma_{\text{TCG}}$, which is responsible for automatically generating test cases based on the question. (2) In the second step, we run MCTS on the original code dataset to generate code dataset with reasoning processes $\mathcal{D}_{\text{process}}$, including a validity indicator to distinguish between correct and incorrect reasoning steps. (3) Once we have data that includes the reasoning process, the third step is to fine-tune the policy model $\pi_{\theta}$, training it to behave in a “think before acting” manner. (4) The reasoning process data can also be used to initialize the process reward model (PRM) $\rho_{\text{PRM}}$, which evaluates the quality of reasoning steps. (5) The fifth step is the most crucial: with PRM $\rho_{\text{PRM}}$ providing process rewards and TCG $\gamma_{\text{TCG}}$ providing outcome rewards, the policy model $\pi_{\theta}$ is updated with reinforcement learning. (6) In the 6th step, based on the updated policy model, new reasoning data can be generated. This new data can then be used to fine-tune the PRM again (4th step). Therefore, steps 4, 5, and 6 form an iterative cycle, where self-play continues to drive model improvements. The flow between the six steps is illustrated in Fig. 2. The following section will introduce each step in detail.
@@ 3.1.3 Experiments
In practical scenarios, the generator’s performance has generally met the requirements for assessing code correctness. Looking ahead, we plan to incorporate the test case generator as an outcome verifier during the inference process. This approach aims to ensure the correctness of generated outputs by validating them against dynamically generated test cases, enabling more robust inference-time search for code generation.
@@ 3.2.1 Pseudocode-based Reasoning Process
The definition of the reasoning process is crucial. As mentioned in the Introduction, we explore a pseudocode-based MCTS approach designed to guide large language models in deep reasoning for complex code tasks. Pseudocode, serving as an intermediate representation between natural language descriptions and actual code, offers a more abstract and concise way to express the logical flow of algorithms or programs. To integrate pseudocode reasoning into step-level Chain-of-Thought (CoT), as illustrated in Fig. 4, we define three key behavioral actions infused with pseudocode reasoning:
@@ 3.4 PRM Training
Point-wise In this format, data collected from the search tree are organized as $D=\{(Q_{i},S_{i}^{1:j-1},S_{i}^{j},v_{i}^{j})\mid i=1,2,\ldots,N\}$, where $N$ is the number of samples, and $v_{i}^{j}$ represents the value label assigned to step $S_{i}^{j}$ during the tree search process. Depending on the processing method, this label can be used to derive either hard or soft estimates. Following the approach in (Wang et al., 2024c), the PRM is trained using the objective:
@@ 4.1 Bitter Lesson: Data is All You Need
Over the last decade, the AI field has been developing along a central line towards maximizing computation-intelligence conversion efficiency (Chung, 2024), which is to efficiently convert the ever-increasing computing power into higher intelligence levels. Along this line, as illustrated at the top of Fig. 6, early advancements prioritized improvements on the model side: from SVM to DNN and then to Transformer, scalable model architectures were designed to fully leverage computational power.
In recent years, the focus has shifted towards the data side. Techniques such as Semi-Supervised Learning (SSL) in pre-training and Reinforcement Learning (RL) in post-training have aimed to harness natural and synthesized data more effectively. The o1 model continues this line. It moves from SFT, which leverages high-quality supervised data, to RLHF, which utilizes environmental feedback to access theoretically unlimited data, and finally to o1’s innovative approach of supervising the generation process through reward signals derived from the generated reasoning process itself.
This progression suggests that, with Transformer architectures now capable of scaling to handle vast amounts of data and training models of sufficient size, the only remaining challenge converges to acquiring adequate data. One approach is to collect data wherever it is lacking, such as reasoning data for system-2 abilities or physical world trajectories for embodied intelligence. Another approach is to explore data types that do not yet exist in the human world, which requires further exploration of techniques like RL and Self-Play.
@@ 4.2 Sweet Lesson: Beyond Human Data
A common criticism of LLM is its reliance on existing human-recorded data, which inherently limits its potential. As Wittgenstein stated, “The limits of my language mean the limits of my world.” The finite scope and depth of human language records constrain the cognitive capabilities of LLMs. However, the success of o1 demonstrates that we can now explore the underlying thought processes behind these recorded data through RL. This advancement signifies a pivotal shift in AI development, moving from mere imitation of human language to the autonomous generation of novel cognitive processes.
More interestingly, these thought process data do not necessarily be confined to natural language. As highlighted in a recent Nature paper, “language serves primarily as a tool for communication rather than the essence of thought.” (Fedorenko et al., 2024) In our observations, some of the thought chains generated by o1 contain nonsensical text, suggesting that the thinking tokens may not correspond to discrete natural language words. If the model has developed itself a more efficient form of internal representation for thinking, this will significantly elevate the efficiency of thought processes and problem-solving mechanisms, not only transcending the limitations imposed by human language data but also further unlocking the potential of model capabilities.
4.3 Opportunities: System1 + X to System2 + X
The self-play RL framework provides a viable solution for exploring underlying data, which opens up the possibility of exploring System-2 solutions for many tasks that were previously reliant on System 1 capabilities. By integrating more thoughtful, step-by-step processes into task execution, we believe that this approach can yield positive results across a wide range of domains (Kant et al., 2024; Ganapini et al., 2021; Valmeekam et al., 2024; Lowe, 2024). Tasks traditionally solved using System 1 capabilities, such as reward modeling (Mahan et al., 2024), machine translation (Zhao et al., 2024), retrieval-augmented generation (RAG) (Li et al., 2024), and multimodal QA (Islam et al., 2024), have already benefited from the deeper reasoning capabilities enabled by System-2 thinking.
The o1 model’s system card demonstrates notable improvements in model safety. Inspired by this, we have recently explored the concept of System-2 Alignment, which involves guiding models to thoroughly evaluate inputs, consider potential risks, and correct biases in their reasoning (Wang & Sang, 2024). We introduced three methods to realize System-2 alignment: prompting, supervised fine-tuning, and reinforcement learning with process supervision. We are applying the Self-Play+RL framework presented in this report to System-2 alignment, aiming to further enhance the model’s ability to think deliberately and reduce vulnerabilities in complex scenarios.
4.4 Challenges: World Model Encoding
@@
For o1-like inference models to be deployed across broader real-world applications, two major challenges need to be addressed, both involving with the RL environments. The first challenge concerns reward function generalization. This has been already discussed in the community. For example, leveraging the enhanced ability of inference models to understand high-level natural instructions, approaches like Constitutional AI (Bai et al., 2022) might directly define reward functions in natural language. Alternative strategy focuses on improving coding capability and transforming the other tasks into coding problems for resolution.
@@
Another less mentioned challenge concerns environment state update during planning. Unlike classic model-free RL methods without planning, such as Q-learning, where state transitions are not explicitly modeled, o1-like planning models rely on behavior simulation and forward search, requiring knowledge of the updated state following an action. This shifts the paradigm towards model-based RL. Fortunately, in well-defined tasks such as programming, mathematics, and Go, the environment dynamics are often deterministic. For example, the world models of Go and other board games can be described explicitly through rules. For programming and mathematics, large language models inherently embed their world models regarding programming syntax and axiomatic logic. These deterministic environment dynamics allow precise computation of state transition probabilities following specific actions.
@@
However, in many real-world applications, such as device use (Wang et al., 2024b; a) and embodied agents, obtaining state updates requires interaction with external environments or simulators. This introduces significant computational and time costs. For example, in device use, behaviors like clicking, inputting, or scrolling must be simulated in a way that involves page rendering, state updates, and sometimes complex backend interactions like network requests. Moreover, o1-like models face the limitation of not being able to perform online behavior simulation during inference, which prevents the model from validating or correcting its actions by returning to a previous state. This leads to inability to backtrack and refine decisions.
@@
Therefore, one of the key directions is to attempt explicit modeling of the environment by developing a world model for state transition prediction. The world model takes as input the current and past states as well as actions, and produces the next state as output. This allows the agent to interact with its internal world model, rather than directly with the real environment or a simulator. However, since accurately building such world models is very difficult, world models have typically been applied to environments where the dynamics are relatively simple and well-understood. The good news is, the recent rapid advancements in interactive content generation (Parker-Holder et al., 2024) and generative games (Sang, 2024) offer promising progress that could facilitate more accurate and practical environment modeling for planning-based reasoning in real-world applications.
@@
Prospects.   The o1 model is clearly influenced by AlphaGo: AlphaGo utilized imitation learning to initialize the policy network, reinforcement learning to fine-tune the policy and learn the value network, and MCTS as an online search strategy, which parallels LLM’s pre-training, post-training, and inference. AlphaGoZero took a more advanced approach by not relying on historical data, which exactly mirrors current trends in LLM development increasingly emphasizing the post-training stage. If we follow their subsequent evolution, we can anticipate similar developments in o1-like reasoning models.
After AlphaGoZero, the Alpha series first developed towards generalization: AlphaZero was applied to Go, Chess, and Shogi. To further tackle more complex scenarios in Atari video games, MuZero requires a dedicated model to handle state transitions. Its approach involves simultaneously updating a world model and a reward model, enabling model-based planning in a latent space rather than relying on explicit environmental observations. In analogy, selecting a compact state representation and constructing an effective world model to support efficient planning are key to applying o1-like models in real-world scenarios for solving long-horizon reasoning tasks. Interestingly, just a week after the release of o1, 1X, the robotics company backed by OpenAI, unveiled its world model project. This initiative aims to develop a predictive framework to simulate and anticipate the outcomes of actions in real-world environments. It highly envisions the potential applications of o1-like reasoning models in advancing embodied intelligence.